Fermi National Laboratory


Accelerator Update for June 6, 2005 - June 8, 2005

Monday June 6
The day shift began with Operations monitoring store 4194, with the Antiproton Source (Pbar) stacking, and with the MiniBooNE and NuMI experiments taking beam. The Meson MCenter experiment is off for a week due to beamline upgrades.

Operations put the Recycler into shot setup to stash antiprotons at 9:57 AM. The stash ended and Pbar resumed stacking at 11:07 AM.

Pbar experts began Debuncher/Accumulator studies at 12:27 PM.

At 4:30 PM, the NuMI Horn tripped off causing the beam permit to drop. Operators reset everything and had NuMI back up six minutes later.

At 5 PM, the NuMI beam permit went down again due to a Horn trip. Operators contacted both power supply and NuMI experts. NuMI beam was disabled.

The Pbar studies ended at 6:31 PM.

Operations resumed sending beam to NuMI at 7:03 PM while experts investigated the Horn problems. At 8:51 PM, experts reported finding and fixing a loose capacitor. NuMI was back on line.

At 9:19 PM, Operations reported sparking in the I- Source. Operators investigated and found a problem with its voltage control module. Operators had the module replaced by10 PM. Pbar resumed stacking.

Tuesday June 7
The midnight shift began with Operations monitoring store 4194 and the stash, with Pbar stacking, and with MiniBooNE and NuMI taking beam.

Operations terminated store 4194 at 3:01 AM and then put the Tevatron (TeV) into shot setup.

Around 2:29 AM, Operations reported a severe drop in the Pbar stack rate. Operators investigated but found nothing wrong, and then at 3:30 AM everything returned to normal. They guessed it might have been due to a dead spot on the target.

Operations established combination store 4195 at 5:21 AM with an initial luminosity of 52.22E30. Operations reported that many protons and antiprotons were lost during the ramp and squeeze due to a corrector magnet set to a wrong value.

Operators prepared the Booster for access at 7:21 AM. After racking out, the operators noticed that a ceramic insulator for GMPS4 was cracked. Operators notified experts.

NuMI was ready for access by 7:36 AM.

Operator prepared the Main Injector (MI) for access at 7:45 AM.

At 12:31 PM, Operations reported that the power was turned off for the Booster East area to allow technicians to repair the broken GMPS insulator. They completed their repairs and had the power back on by 1:54 PM.

Operations reported at 5:01 PM that MI was on and ramping.

Operations reported at 6:09 PM that there was beam in the Booster.

At 6:29 PM, Operations reported that the NuMI safety system wouldn't make up. An operator contacted an interlock technician.

Operations established beam to the MI dump at 6:33 PM.

Pbar began stacking at 6:40 PM.

Operations established beam to MiniBooNE at 7:06 PM.

At 9:31 PM, Operations reported that the TeV sector B air compressor tripped off. Operators investigated and managed to get it back on by 9:58 PM.

The interlock tech finished his NuMI repairs and Operations established beam to the experiment at 11:46 PM.

Wednesday June 8
The midnight shift began with Operations monitoring store 4195, with Pbar stacking, and with MiniBooNE and NuMI taking beam.

At 1:19 AM, Operations reported a problem with the NuMI BPMs. An operator called in an expert who had the problem resolved by 2:56 AM.

Around 1:19 AM, Operations reported that Pbar began unstacking antiprotons due to an emittance blowup. Operators completed a tune change around 1:49 AM and normal stacking resumed.

Operations put the Recycler into shot setup to stash antiprotons at 2:03 AM. The stash was completed by 2:56 AM. Stacking resumed a few minute later.

At 3:58 AM, a Booster kicker (MKS04) tripped off. Operators investigated and had the kicker back on six minutes later.

At 5:52 AM, a Booster power supply (GMPS) tripped off. Operations called in an expert. This disabled the Booster beam.

TeV experts began an end-of-store study at 6:26 AM.

Operations prepared the Booster for an access at 6:43 AM.

Operations terminated store 4195 at 6:52 AM and began preparing the Collider Detector at Fermilab (CDF) and D-Zero (D0) for access.

The Booster experts reported at 7:04 AM that GMPS had tripped due to another bad capacitor. They jumpered out the bad capacitor; the power supply will be repaired later. Booster beam resumed at 7:45 AM.

The Plan for Wednesday
The plans for today are to allow CDF and D0 to complete the collision hall work and then let TeV experts conduct a study. Operations will establish a new store this afternoon.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/8/2005   email Fermilab